OpenStudy (anonymous):

Solve: −2(2x + 5) − 3 = −3(x − 1) -4x+10-3=3x-3 x+10-3=3 at this point do i cancel out 3's or do I add everthing manking the answer 16

OpenStudy (anonymous):

-4x-10-3=-3x+3 -13-3=-3x+4x x=-16

OpenStudy (anonymous):

−2(2x + 5) − 3 = −3(x − 1) -4x-10-3=-3x+3 4x+13=3x-3 x=-16

OpenStudy (anonymous):

Great, thanks

OpenStudy (anonymous):

- 2( 2x +5)-3= -3(x-1) Solving we get -4x-10-3= -3x +3 -16=x

OpenStudy (anonymous):

thanks nishika17